A bedroom with a wall area of 6912 square meters is to be covered by wallpaper. If the size of a roll of wallpaper is 16 meters long and 48 meters wide, how many rolls of wallpaper is needed to cover this area?
1
Expert's answer
2012-08-10T10:38:21-0400
The area of a roll of wallpaper is

A = 16*48 = 768m²,

so there needed

N = 6912/768 = 9

rolls to cover the walls.